Mark is 8 years younger than Gerry. In ten years, he will be ten years more than one-half as old as Gerry. Find their present ages.
Answer by stanbon(75887)

Mark is 8 years younger than Gerry. In ten years, he will be ten years more than one-half as old as Gerry. Find their present ages.
---------
Equations:
m = g - 8
m+10 = (1/2)(g+10)+10
------
Substitute for "m" and solve for "g":
g-8+10 = (1/2)(g+10)+10
------
g - 8 = (1/2)(g+10)
2g - 16 = g+10
g = 26 (Gerry's age now)
----
Solve for "m":
m = g - 8
m = 24 years (Mark's age now)
-----
Cheers,
Stan H.
===========